OpenStudy (anonymous):

Help please: (1-sqrt(3)*i)^5 using De Moivre's theorem

OpenStudy (anonymous):

\[(1-\sqrt{3}i)^{5}\]

OpenStudy (anonymous):

I think \(r=2\): \((\large1-\sqrt3 i)^5=[2(\frac{1}{2}-\frac{\sqrt3}{2}i)]^5=[2(cos\frac{\pi}{3}-isin\frac{\pi}{3})]^5\)

OpenStudy (anonymous):

How did you get the 1/2 and \[\sqrt{3}/2\]

OpenStudy (anonymous):

Take 2 outside: \(\large1+\sqrt3=2(\frac{1}{2}+\frac{\sqrt3}{2})\)

OpenStudy (anonymous):

I mean -. Sorry

OpenStudy (anonymous):

Why is r=2?

OpenStudy (anonymous):

and how did you find \[\theta \]?

OpenStudy (anonymous):

This is Moirve's theorem: \([r(cos~\theta + isin~\theta)]^n=r^n(cos~n\theta +isin~n\theta)\), right? So you know why r = 2. Now we have to find \(\theta\), knowing that \(cos~\theta=\frac{1}{2}\) and \(sin~\theta=\frac {\sqrt3}{2}\). Then \(\theta=\frac{\pi}{3}\)

OpenStudy (anonymous):

so what would be the simplified answer?

OpenStudy (anonymous):

So \([2(cos\frac{\pi}{3}+isin\frac{\pi}{3})]^5=2^5(cos\frac{5\pi}{3}+isin\frac{5\pi}{3})=...\) Can you simplify that?

OpenStudy (anonymous):

\[16-16\sqrt{3}i\]

OpenStudy (anonymous):

But it's wrong

OpenStudy (anonymous):

Oops. I am so sorry. It should be \(2^5(cos\frac{5\pi}{3}-isin\frac{5\pi}{3})\). My fault

OpenStudy (anonymous):

Okay thank you so much!

OpenStudy (anonymous):

You are welcome